40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Feb 22 07:53:06 2022, Yehonathan, Update, WOPO, Waking up WOPO 
    Reply  Fri Feb 25 17:09:53 2022, shruti, Update, WOPO, Waking up WOPO - green beam 
       Reply  Wed Mar 2 14:38:55 2022, Yehonathan, Update, WOPO, Waking up WOPO - some more fiddling and a plan 
          Reply  Fri Mar 4 11:49:45 2022, shruti, Update, WOPO, Waking up WOPO - attempts at readout LO_shot_noise.pdf
             Reply  Thu Mar 10 17:12:01 2022, awade, Update, WOPO, Waking up WOPO - attempts at readout 
                Reply  Wed Mar 16 12:22:44 2022, Yehonathan, Update, WOPO, Waking up WOPO - attempts at readout Moku_Noise.pdfPD_Connected_no_light.pdfDiff_Channel.pdf
                   Reply  Tue Mar 22 09:19:38 2022, shruti, Update, WOPO, Waking up WOPO - back to green path before_opt.jpgaround_waveguide.png
Message ID: 2735     Entry time: Tue Mar 22 09:19:38 2022     In reply to: 2733
Author: shruti 
Type: Update 
Category: WOPO 
Subject: Waking up WOPO - back to green path 

[Yehonathan, Shruti]

Yesterday we went back to fiddling with the green path. Soon after opening the green shutter and then switching the doubling cavity to 'AUTO' we were able to see 150 mW of green light. We were able to replicate this a couple of times yesterday.

Since we had earlier removed the green fiber from the fiber launch to clean its tip, the coupling into the fiber turned out to be quite poor. As can be seen in Attachment 1, Yehonathan pointed out that a lot of green light was being lost to the cladding due to poor coupling. He then played around with the alignment and finally was able to see 65% coupling efficiency. This process seemed to involve a great amount of trial and error through several local power minima.

Attachment 2 shows that the coupling between the two fibers at the 532 nm input of the waveguide is quite poor (there is visible light being lost in the cladding). Furthermore, this light intensity decreases as we get closer to the waveguide meaning this light is being dissipated in the fiber. Even at the 1064 nm output where we expect to see squeezing there is some remnant green light.

We wanted to test if the green leakage reaching the PDs were causing additional noise. For this we just looked at the spectrum analyzer on the Moku (after amplifying 100x with the SR 560) and saw no difference in the noise spectrum with and without the green shutter being open. Although, we're not convinced with this measurement since we were not able to find good quality SMA cables for the entire path. Moving around the BNCs seemed to change the noise. Also, near the end, we noticed some coupling between the two channels on the Moku while measuring the noise that seemed to cause additional noise in one of the channels. We did not have sufficient time yesterday to probe this further.

Attachment 1: before_opt.jpg  1.031 MB  Uploaded Tue Mar 22 10:21:01 2022  | Hide | Hide all
before_opt.jpg
Attachment 2: around_waveguide.png  3.970 MB  Uploaded Tue Mar 22 10:22:27 2022  | Hide | Hide all
around_waveguide.png
ELOG V3.1.3-